Find the value dy/dx if y=1+ sinx - cosx ÷1+ sinx + cosx

Grade:12

2 Answers

Ananya
6 years ago
Y=1+sinx-cosx+sinx+cosxY=1+2sinxtaking log on both sides logy=log1+log2sinx1/y.dy/dx=1/2sinx.2cosxdy/dx=y(cotx)dy/dx=[1+2sinx]cotx
urvashi verma
6 years ago
firstly,try to make y in a simpler form by using formulas as [1-cosx = 2(sin2x/2) ],[ 1+cosx =2(cos2x/2) ] and sinx = 2 sin(x/2) .cos(x/2).after putting these in your question and taking 2sin(x/2) common from numerator and 2 cos(x/2) from denominator your Y will reduced to tan (x/2) and thus its derivative is given by sec2(x/2) / 2.